# SQL-GEN: Bridging the Dialect Gap for Text-to-SQL Via Synthetic Data And Model Merging

### v1

@@ 3 Experiment
To evaluate the quality of our generated queries, we used benchmark datasets tailored for three dialects. For SQLite, we use two datasets from BIRD: 1) the development set and 2) the mini development set. For PostgreSQL, we utilize three benchmarks: 1) BIRD queries transpiled to PostgreSQL, 2) BIRD PostgreSQL mini development set, and 3) Pagila—a dataset specific to PostgreSQL containing real-world queries originally written for PostgreSQL, which are extracted from online resources. For BigQuery, we use two datasets: 1) BIRD queries transpiled to BigQuery, and 2) the GitHub_repositories dataset, a public BigQuery dataset featuring BigQuery-specific sample question/SQL pairs obtained from tutorials and online resources. Further details of the datasets are provided in Section A.2.
@@
In order to evaluate the quality of the synthetic queries generated with SQL-GEN, we compare the performance of the models trained on synthetic data with the following datasets: 1) Gretel (Gretel, 2024): Gretel Text-to-SQL dataset consists $>$100K high-quality synthetic Text-to-SQL samples with a coverage across 100 distinct domains. 2) SQL create context (b mc2, 2023): This dataset consists $\sim$78K samples, obtained from the Spider (Yu et al., 2018b) and WikiSQL (Zhong et al., 2017) datasets by cleaning these sources. All queries were generated through a human-in-the-loop process, making it a strong baseline for comparing LLM-generated data with human-annotated data. 3) BIRD train set (Li et al., 2024b): This dataset consists of $\sim$10k human-annotated samples. Queries are considered more complex in comparison to the Spider and WikiSQL benchmarks.
For dialects other than SQLite, all queries from the baselines are transpiled to the target dialects to ensure their validity. Details of the models and metrics are provided in Section A.3 and details of the seed templates and tutorials are provided in Section A.4. Additionally, in Section A.7, we provided the analysis on dialect-specific keywords across different datasets.
@@ 3.1 SQLite Results
SQL-GEN generated samples significantly surpass the Gretel Text-to-SQL dataset, achieving a large performance gain of approximately 10% across all model sizes. Furthermore, LLMs trained on SQL-GEN synthetic data consistently outperform those trained on the human-annotated SQL Create Context data, underscoring the high quality of SQL-GEN synthetic data. While LLMs trained on the BIRD dataset consistently exhibit the highest performance on BIRD development sets, this outcome is likely due to overfitting to the canonical input distribution of the BIRD train set which is similar to its development set (Yu et al., 2020).
Table 1: Execution Accuracy (EX) of SQLite Models on the BIRD development set and minidev set using CodeLlama 7B, CodeGemma 7B, and Codestral 22B Models. "-" denotes the zero-shot performance of the models. OS refers to the dataset that is created using open-source LLMs (Llama 3 + Mixtral)
@@
Our synthetic dataset (OS) CodeLlama 7B dev set 34.55 +16.69 minidev set 28.4 +14.8
Our synthetic dataset (Gemini) CodeLlama 7B dev set 38.33 +20.47 minidev set 30.00 +16.4
@@
Our synthetic dataset (OS) CodeGemma 7B dev set 41.13 +7.37 minidev set 35.6 +5.6
Our synthetic dataset (Gemini) CodeGemma 7B dev set 42.37 +8.64 minidev set 36.4 +6.4
@@
Our synthetic dataset (OS) Codestral 22B dev set 47.52 +3.0 minidev set 44.4 +4.0
Our synthetic dataset (Gemini) Codestral 22B dev set 50.45 +5.93 minidev set 46.6 +6.2
@@
To pinpoint the gain is only for simple SQL queries or not, we evaluate different models in terms of their performance on different SQL query complexities of simple, medium, and challenging, presented in Section A.5.
3.1.1 Database Adaptation
SQL-GEN operates independently of specific databases, enabling the generation of high-quality synthetic data for any database. Therefore, as another use case of synthetic data, we introduce Database Adaptation, to improve the performance for cross-domain Text-to-SQL setting. This involves generating synthetic queries for databases for which no pre-existing question-SQL pairs are available. We apply this training in two distinct ways: (1) in-context learning, which leverages the generated queries directly within the model’s input context, and (2) model tuning, which involves supervised fine-tuning of the model weights:
Database Adaptation With Model Tuning:
Our synthetic data generation pipeline is designed to generate question-SQL pairs for any database. To demonstrate this, we generated 10K pure synthetic question-SQL pairs across the 11 databases in the BIRD development set and separately 10K samples for the entire databases in the BIRD training set using Llama 3 and Mixtral. We then compared the performance of two model against a model trained on the original 10K training samples from the BIRD benchmark.The results are detailed in Table 4. The table indicates that our synthetic generation approach on the BIRD development set databases achieves performance comparable to the original BIRD training set with only 1.5% gap. This is particularly noteworthy given that generating synthetic samples is significantly less resource- and cost-intensive compared to creating 10,000 human-annotated samples. The latter involves 11 crowdsource workers to annotate the samples. Additionally, synthetic data generation on development split outperforms train split showing that SQL-Gen helps to learn unseen database and improve performance.
@@ 3.1.1 Database Adaptation
Table 4: Evaluating the database adaptation with CodeLlama 7B model on BIRD dev set.
Database Adaptation with In-context Learning:
@@
3.1.2 Data Augmentation
Beyond merely creating a pool of pure synthetic question-SQL pairs for training, synthetic data generation offers the potential to augment existing datasets (e.g. mixing synthetic data with original dataset), thereby enhancing model performance beyond what is achievable with solely the available data. This section delves into the possibilities of integrating synthetic data generated for specific target databases (as discussed in Database Adaptation, see Section 3.1.1) with pre-existing training datasets. To this end, we merge 10K synthetic question-SQL pairs generated on the BIRD development databases with the 10K pairs from the BIRD training set. We then trained various models using this combined dataset and compared their performance to models trained solely on the original BIRD training set. For a balanced comparison, models using the combined datasets are trained for only one epoch, whereas those trained exclusively on the BIRD training set underwent two epochs of training. As shown in Table 5, augmenting training data with our synthetic dataset results in a performance improvement of up to 5.6%. This represents a significant enhancement compared to previous works, such as Yang et al. (2024), which demonstrated only a 1.5% improvement in performance after augmentation on the same base model CodeLLama 7B.
@@ 3.1.2 Data Augmentation
3.2 PostgreSQL Results
Table 6: Execution Accuracy (EX) of PostgreSQL Models on the three PostgreSQL benchmarks using CodeLlama 7B and Codestral 22B. "-" denotes the zero-shot performance of the models. OS refers to using open-source models of Llama3 and Mixtral for data generation.
Our synthetic dataset (OS) PostgreSQL BIRD CodeLlama 7B 38.69 +14.4 Codestral 22B 49.94 +3.36
Our synthetic dataset (Gemini) PostgreSQL BIRD CodeLlama 7B 39.22 +14.93 Codestral 22B 49.84 3.26
- PostgreSQL BIRD CodeLlama 7B 24.29 0 Codestral 22B 46.58 0
Our synthetic dataset (OS) PostgreSQL Minidev CodeLlama 7B 20.8 +7.6 Codestral 22B 33.0 +1.2
Our synthetic dataset (Gemini) PostgreSQL Minidev CodeLlama 7B 25.4 +12.2 Codestral 22B 33.0 +2.2
- PostgreSQL Minidev CodeLlama 7B 13.2 0 Codestral 22B 31.8 0
Our synthetic dataset (OS) Pagila CodeLlama 7B 39.13 +15.22 Codestral 22B 52.27 +2.27
Our synthetic dataset (Gemini) Pagila CodeLlama 7B 39.13 +15.22 Codestral 22B 50 0.0
- Pagila CodeLlama 7B 23.91 0 Codestral 22B 50 0
3.3 BigQuery Results
Similar to the PostgreSQL experiments, we present the results of the CodeLlama 7B and Codestral 22B model trained on various baseline datasets and evaluated on two BigQuery benchmark datasets: BIRD and the GitHub Repository database. Looking at the results provided in Table 7, consistent with the trends observed for the PostgreSQL dialect, on BigQuery BIRD, the model trained on our generated samples achieves the second highest performance, following the BIRD train set. For the GitHub Repository database, which is a BigQuery dialect specific dataset, our model outperforms the second-best model by a 10% margin, further demonstrating the effectiveness of our method to train dialect specific models.
Table 7: Execution Accuracy (EX) of BigQuery Models on the two BigQuery benchmarks using CodeLlama 7B and Codestral 22B. "-" denotes the zero-shot performance of the models.
Our synthetic dataset (OS) BigQuery BIRD CodeLlama 7B 32.08 +15.51 Codestral 22B 45.22 +7.03
Our synthetic dataset (Gemini) BigQuery BIRD CodeLlama 7B 33.53 +16.96 Codestral 22B 47.24 +9.05
- BigQuery BIRD CodeLlama 7B 16.57 0 Codestral 22B 38.19 0
Our synthetic dataset (OS) Github Repository CodeLlama 7B 27.5 +12.5 Codestral 22B 30.0 +10.0
Our synthetic dataset (Gemini) Github Repository CodeLlama 7B 25.0 +10.0 Codestral 22B 30.0 +10.0
- Github Repository CodeLlama 7B 15 0 Codestral 22B 20.0 0
@@ 3.4 Experts Merging Results
We evaluate various expert merging approaches for integrating dialect-specific models into a single unified model and compare to our method based on the Mixture of Experts (MoE) architecture. We utilize three expert CodeLlama 7B models, each trained on synthetic question-SQL pairs for SQLite, PostgreSQL, and BigQuery. We consider three popular model merging techniques: DARE (Yu et al., 2024), TIES (Yadav et al., 2024), and SLERP. Unlike the first two, SLERP can only merge two models at a time. Therefore, we initially merge the SQLite and PostgreSQL experts and then combined the resulting model with the BigQuery expert. Additionally, we fine-tune a generalist (not dialect-specific) CodeLlama 7B and MoE 3x7B model on 40K samples from a mix of different dialects to establish a baseline for comparison. The generalist MoE baseline is constructed from CodeLlama 7B with three experts being MLP layers repeated three times. We train with 40K combined dialect samples for 1 epoch. We compare all these methods to the proposed method for initializing the MoE model where we merge the self-attention sub-layers using the SLERP method and created three distinct MLP experts from the dialect experts’ MLP layers, training only for one epoch of 20K samples from different dialects. The MoE routers are initialized using hidden states from dialect-specific keyword prompts. We train for a single epoch to promote effective collaboration among the submodules. We also included the performance of our proposed MoE model before the single epoch fine-tuning to better understand the effect of coordination through fine-tuning. We assess the performance of the different models on PostgreSQL’s Pagila, BigQuery’s Github Repository, and 10% of random samples from the SQLite BIRD dev set, with results detailed in Table 8. The table demonstrates that our proposed MoE model overall performance outperforms other methods and even exceeds the performance of the individual dialect experts, highlighting the effectiveness of our approach in sharing common SQL knowledge across dialects while preserving dialect-specific expertise. The MoE architecture also enhances the model’s learning capacity, contributing to improved overall performance. Notably, our initialization method is more effective at maintaining high dialect-specific performance compared to the generalist MoE 3x7B model. Among all merging techniques, SLERP achieves the highest performance, surpassing even the generalist model trained on the combined dialect-specific datasets, which is the main reason for initializing the self-attention sub-layers. Moreover, the results suggest that our proposed method for initialization even before fine-tuning provide a strong baseline surpassing TIES and DARE method for model merging. In Section A.9 we provide detailed analysis of token-level routing for MoE meodels.
@@ 4 Conclusions
We present a novel method for generating dialect-specific synthetic data to tackle the diverse SQL dialect challenges in Text-to-SQL systems. This method addresses the unique keywords and functions of each SQL dialect, offering a scalable solution across various dialects. It significantly narrows the performance gap with human-annotated datasets for SQLite and creates the highest quality datasets for other dialects. Our comprehensive evaluations across three models and multiple benchmarks, showcase the effectiveness of the proposed approaches for data generation. Additionally, our innovative approach integrates dialect-specific experts into a unified model, enhancing performance by promoting effective information sharing among them.
@@ A.4.1 Quality Check Ablation
Table 10: Performance Comparison of Two Scenarios: One Where the Same Model Generates and Filters Candidate Samples, and Another Where a Secondary Model Is Used for Filtering.
@@ A.6.4 Deduplication And Length Check
A.7 SQL keywords
A.8 Dialect Specific Keywords
@@
A.9 MoE Analysis
In this section, we analyze the hidden representations of our proposed MoE model and compare it with the baseline generalist MoE model across three distinct layers: Layer 1, Layer 16, and Layer 32. Although both MoE models are trained with a load balancing loss, our initialization approach for the gates leads to an expert collapse in the middle layers. This issue primarily stems from the high similarity in the hidden representations of the positive prompts for each dialect expert across all layers. Additionally, similar to the experiments conducted with the original Mixtral model (Jiang et al., 2024), there is no distinct expert associated with different token types in either our MoE model or the baseline generalist MoE model.

### v2

@@ 3 Experiments
We use benchmark datasets tailored for three dialects. For SQLite, we use two datasets from BIRD: 1) the development set and 2) the mini development set. For PostgreSQL, we utilize three benchmarks: 1) BIRD queries transpiled to PostgreSQL, 2) BIRD PostgreSQL mini development set, and 3) Pagila—a dataset specific to PostgreSQL containing real-world queries originally written for PostgreSQL, which are extracted from online resources. For BigQuery, we use two datasets: 1) BIRD queries transpiled to BigQuery, and 2) the GitHub_repositories dataset, a public BigQuery dataset featuring BigQuery-specific sample question/SQL pairs obtained from tutorials and online resources. Further details of the datasets are provided in Section A.2.
@@
In order to evaluate the quality of the synthetic queries generated with SQL-GEN, we compare the performance of the models trained on synthetic data considering the following datasets: 1) Gretel (Gretel, 2024): Gretel Text-to-SQL dataset consists $>$100K high-quality synthetic Text-to-SQL samples with a coverage across 100 distinct domains. 2) SQL create context (b mc2, 2023): This dataset consists $\sim$78K samples, obtained from the Spider (Yu et al., 2018b) and WikiSQL (Zhong et al., 2017) datasets by cleaning these sources. All queries were created using a human-in-the-loop approach, providing a robust baseline for comparing data generated by LLMs to human-annotated data. 3) BIRD train set (Li et al., 2024b): This dataset consists of $\sim$10k human-annotated samples. Queries are considered more complex in comparison to the Spider and WikiSQL benchmarks.
For dialects other than SQLite, all queries from the baselines are transpiled to the target dialects to ensure their validity. Details of the models and metrics are provided in Section A.3 and details of the seed templates and tutorials are provided in Section A.4.
3.1 Dialect-specific SQL keywords converge improvement
3.2 PostgreSQL Results
Table 1: Execution Accuracy (EX) of PostgreSQL Models on the three PostgreSQL benchmarks using CodeLlama 7B and Codestral 22B. "-" denotes the zero-shot performance of the models.
Our synthetic dataset PostgreSQL BIRD CodeLlama 7B 39.22 +14.93 Codestral 22B 49.84 3.26
- PostgreSQL BIRD CodeLlama 7B 24.29 0 Codestral 22B 46.58 0
Our synthetic dataset PostgreSQL Minidev CodeLlama 7B 25.4 +12.2 Codestral 22B 33.0 +2.2
- PostgreSQL Minidev CodeLlama 7B 13.2 0 Codestral 22B 31.8 0
Our synthetic dataset Pagila CodeLlama 7B 39.13 +15.22 Codestral 22B 50 0.0
- Pagila CodeLlama 7B 23.91 0 Codestral 22B 50 0
3.3 BigQuery Results
Similar to the PostgreSQL experiments, we present the results of the CodeLlama 7B and Codestral 22B model trained on various baseline datasets and evaluated on two BigQuery benchmark datasets: BIRD and the GitHub Repository database. Looking at the results provided in Table 2, consistent with the trends observed for the PostgreSQL dialect, on BigQuery BIRD, the model trained on our generated samples achieves the second highest performance, following the BIRD train set. For the GitHub Repository database, which is a BigQuery dialect-specific dataset, our model outperforms the second-best model by a 10% margin, further demonstrating the effectiveness of our method to train dialect specific models.
Table 2: Execution Accuracy (EX) of BigQuery-specific models on the two BigQuery benchmarks using CodeLlama 7B and Codestral 22B. "-" denotes the zero-shot performance of the models.
Our synthetic dataset BigQuery BIRD CodeLlama 7B 33.53 +16.96 Codestral 22B 47.24 +9.05
- BigQuery BIRD CodeLlama 7B 16.57 0 Codestral 22B 38.19 0
Our synthetic dataset Github Repository CodeLlama 7B 25.0 +10.0 Codestral 22B 30.0 +10.0
- Github Repository CodeLlama 7B 15 0 Codestral 22B 20.0 0
@@ 3.4 SQLite Results
SQL-GEN generated samples significantly surpass the Gretel dataset, achieving a large performance gain of approximately 10% across all model sizes. Furthermore, LLMs trained on SQL-GEN synthetic data consistently outperform those trained on the human-annotated SQL Create Context data, underscoring the high quality and utility of the SQL-GEN synthetic data. While LLMs trained on the BIRD dataset consistently exhibit the highest performance on BIRD development sets, this outcome is likely due to overfitting to the canonical input distribution of the BIRD train set which is similar to its development set (Yu et al., 2020).
Table 3: Execution Accuracy (EX) of SQLite Models on the BIRD development set and minidev set using CodeLlama 7B, CodeGemma 7B, and Codestral 22B Models. "-" denotes the zero-shot performance of the models.
@@
Our synthetic dataset CodeLlama 7B dev set 38.33 +20.47 minidev set 30.00 +16.4
@@
Our synthetic dataset CodeGemma 7B dev set 42.37 +8.64 minidev set 36.4 +6.4
@@
Our synthetic dataset Codestral 22B dev set 50.45 +5.93 minidev set 46.6 +6.2
@@
To pinpoint whether the gains are consistent across, we evaluate different models on different SQL query complexity levels: simple, medium, or challenging, as presented in Section A.5.
Database Adaptation:
SQL-GEN operates independently of specific databases, enabling the generation of high-quality synthetic data for any database. Therefore, as another use case of synthetic data, we introduce Database Adaptation, to improve the performance in cross-domain Text-to-SQL setting. This involves generating synthetic queries for databases for which no pre-existing question-SQL pairs are available. We apply this training in two distinct ways: (1) in-context learning, which leverages the generated queries directly within the model’s input context as demonstrations, and (2) model tuning, which involves supervised fine-tuning of the model weights:
Adaptation With Model Tuning:
Our synthetic data generation pipeline is designed to generate question-SQL pairs for any database. To demonstrate this, we generated 10K pure synthetic question-SQL pairs across the 11 databases in the BIRD development set and separately 10K samples for the entire databases in the BIRD training set using Gemini-1.5-pro. We then compared the performance of two model against a model trained on the original 10K training samples from the BIRD benchmark. The results are detailed in Table 5. The table indicates that our synthetic generation approach on the BIRD development set databases achieves performance comparable to the original BIRD training set with only 1.5% gap. This is particularly noteworthy given that generating synthetic samples is significantly less resource- and cost-intensive compared to creating 10,000 human-annotated samples. The latter involves 11 crowd-source workers to annotate the samples. Additionally, synthetic data generation on development split outperforms train split showing that SQL-Gen helps to learn unseen database and improve performance.
@@
Adaptation with In-context Learning:
@@
Data Augmentation:
Beyond merely creating a pool of pure synthetic question-SQL pairs for training, synthetic data generation offers the potential to augment existing datasets (e.g. mixing with original dataset), thereby enhancing model performance beyond what is achievable with solely the available data. We consider integrating synthetic data generated for specific target databases (as discussed in Database Adaptation, see Section 3.4) with pre-existing training datasets. To this end, we merge 10K synthetic question-SQL pairs generated on the BIRD development databases with the 10K pairs from the BIRD training set. We then train various models using this combined dataset and compared their performance to models trained solely on the original BIRD training set. For a balanced comparison, models using the combined datasets are trained for only one epoch, whereas those trained exclusively on the BIRD training set are trained fro two epochs. As shown in Table 6, augmenting training data results in a performance improvement of up to 5.6%, a significant enhancement compared to previous work, such as Yang et al. (2024) (which demonstrated only a 1.5% improvement in performance after augmentation on the same base model CodeLLama 7B).
@@ 3.5 Experts Merging Results
We evaluate various expert merging approaches for integrating dialect-specific models into a single unified model and compare to the proposed approach based on the Mixture of Experts (MoE) architecture. We utilize three expert CodeLlama 7B models, each trained on synthetic question-SQL pairs for SQLite, PostgreSQL, and BigQuery. We consider three popular model merging techniques: DARE (Yu et al., 2024), TIES (Yadav et al., 2024), and SLERP. Unlike the first two, SLERP can only merge two models at a time. Therefore, we initially merge the SQLite and PostgreSQL experts and then combined the resulting model with the BigQuery expert. Additionally, we fine-tune a generalist (not dialect-specific) CodeLlama 7B and MoE 3x7B model on 40K samples from a mix of different dialects to establish a baseline for comparison. The generalist MoE baseline is a MoE 3x7B model initialized from CodeLlama 7B model and trained with 40K combined dialect samples for 1 epoch. We compare all these methods to the proposed method for initializing the MoE model which is trained only for one epoch of 20K samples from different dialects. We train for a single epoch to promote effective collaboration among the submodules. We also include the performance of the proposed MoE model before the single epoch fine-tuning to better understand understand the effectiveness of the proposed initialization method. We assess the performance of the different models on PostgreSQL’s Pagila, BigQuery’s Github Repository, and 10% of random samples from the SQLite BIRD dev set, with results detailed in Table 7. The table shows that the proposed MoE model outperforms other models, even surpassing the individual dialect experts. This highlights the effectiveness of our approach in sharing common SQL knowledge across dialects while maintaining dialect-specific expertise. The MoE architecture also enhances the model’s learning capacity, contributing to improved overall performance. Notably, our initialization method is more effective at maintaining high dialect-specific performance compared to the generalist MoE 3x7B model. Among all merging techniques, SLERP achieves the highest performance, surpassing even the generalist model trained on the combined dialect-specific datasets, which is the main reason for initializing the self-attention sub-layers. Moreover, the results suggest that our proposed method for initialization even before fine-tuning provides a strong baseline, surpassing TIES and DARE methods for model merging. In Section A.8, we provide detailed analysis of token-level routing for MoE architecture.
@@ 4 Conclusions
This work addressed the challenge of cross-dialect generalization in Text-to-SQL by introducing SQL-GEN, a novel framework for generating high-quality synthetic training data from dialect-specific tutorials. The proposed framework addresses the unique challenges such as keywords and functions being different for each SQL dialect, constituting a scalable approach. It significantly narrows the performance gap with human-annotated datasets and creates the highest quality datasets for other dialects. Our comprehensive evaluations across three models and multiple benchmarks showcase the effectiveness of the proposed data generation framework. To further enhance cross-dialect performance, we proposed a novel Mixture-of-Experts (MoE) initialization method that leverages shared knowledge across dialects. By merging self-attention layers from dialect-specific models and initializing expert gates with dialect-specific keywords, we created a unified model capable of outperforming single-dialect models. Our findings demonstrate the effectiveness of SQL-GEN in generating high-quality synthetic data for diverse SQL dialects and highlight the potential of our MoE initialization method for building robust, multi-dialect Text-to-SQL systems. This work opens avenues for future research in cross-lingual Text-to-SQL and the application of synthetic data generation to other code generation tasks.
@@ A.1.1 Synthetic Data Generation
A.1.2 Model Merging
Training specialized, task-specific models presents several challenges, including the storage costs associated with maintaining multiple models, the substantial memory requirements for deploying these models, and the rapid obsolescence of models as training datasets age. One proposed solution to mitigate these issues is model merging (Goddard et al., 2024). Initial approaches to model merging, such as Task Arithmetic (Ilharco et al., 2022), involve calculating task-specific vectors by determining the weight differences between the fine-tuned model and its base counterpart. These vectors are then linearly combined and reintegrated with the original base model. Subsequent methodologies like DARE, TIES, and Model BreadCrumbs (Yadav et al., 2024; Yu et al., 2024; Davari & Belilovsky, 2023) have aimed to minimize interference among task-specific models through techniques such as sparsification, sign consensus algorithms, and the exclusion of extreme values. Additionally, DARE introduces random pruning to align more closely with the base model’s performance (Goddard et al., 2024). More recently, the integration of model merging with Mixture of Experts (MoE) architectures has been explored. This method, termed FrankenMoEs, initializes MoE MLP layers using weights from task-specific models (Goddard, 2024; Tang et al., 2024). Our work extends these efforts by specifically leveraging features from dialect-specific models for gate initialization and merging self-attention sublayers within transformer architectures.
@@ A.4.1 Quality Check Ablation
Table 9: Performance comparison between two scenarios, when the same model generates and filters candidate samples, and another when a secondary model is used for filtering.
@@ A.6.4 Deduplication And Length Check
A.7 Dialect Specific Keywords
@@
A.8 MoE Analysis
We analyze the hidden representations of our proposed MoE model and compare it with the baseline generalist MoE model across three distinct layers: Layer 1, Layer 16, and Layer 32. Although both MoE models are trained with a load balancing loss, our initialization approach for the gates leads to an expert collapse in the middle layers. This issue primarily stems from the high similarity in the hidden representations of the positive prompts for each dialect expert across all layers. Additionally, similar to the experiments conducted with the original Mixtral model (Jiang et al., 2024), there is no distinct expert associated with different token types in either our MoE model or the baseline generalist MoE model.